OpenStudy (anonymous):

A line passes through (3, –2) and (6, 2). Write an equation for the line in point-slope form. Rewrite the equation in standard form using integers.

OpenStudy (anonymous):

@mathisblockingmypath if i give you this formula do you know how to solve it? y-y1=m(x-x1) and the slope is 4/3..

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i just tried

OpenStudy (anonymous):

slope would be the m right ??

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

okay let me try and plug this is

OpenStudy (anonymous):

y-y1=4/3 (x-x1) right @brianayala

OpenStudy (anonymous):

yeah.. let me help you.. y-2= 4/3(x-3)

OpenStudy (anonymous):

the slope is one and one third

OpenStudy (anonymous):